OpenStudy (anonymous):

Tony has 6 dollars in change in his car. Every time he drives to visit his grandma, he pays $1.25 in tolls. Which graph represents the change Tony has in his car? In a graph please and explanation

OpenStudy (anonymous):

@mathstudent55 @iambatman

OpenStudy (anonymous):

@texaschic101 @paki

OpenStudy (texaschic101):

you have any pictures of the graphs ?

OpenStudy (anonymous):

I can take a screenshot if that would help @texaschic101

OpenStudy (texaschic101):

that might help, if you don't mind

OpenStudy (anonymous):

Alright, hold on~

OpenStudy (texaschic101):

that would be graph 3....he is starting at 6.....coming down 1.25 each time he pays the toll

OpenStudy (anonymous):

The answer is C?

OpenStudy (texaschic101):

I don't see a C..but it is the 3rd graph

OpenStudy (anonymous):

Could you explain? ;w; I'm completely confused.. And could I ask another question ?

OpenStudy (texaschic101):

let me check again...hold on

OpenStudy (anonymous):

I think i guessed that the 1st try and I think i got it wrong

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) (0,6)...x1 = 0 and y1 = 6 (8,-4)...x2 = 8 and y2 = -4 now we sub slope = (-4 - 6) / (8 - 0) slope = -10/8 slope = - 5/4 or -1 1/5 or - 1.25 I found it..sorry about that..it is graph D, or the last graph, because the slope is -1.25 and thats what you want

OpenStudy (texaschic101):

what you needed to do was to use the slope formula, and plug in 2 sets of points. And when you work it out, it comes out to -1.25....you want it to be a negative slope because he is paying money, not getting money.

OpenStudy (texaschic101):

Did I just totally confuse the hell out of you ??

OpenStudy (texaschic101):

or are you kind of understanding this

OpenStudy (anonymous):

Wait hold on I'm sorry parents were fighting

OpenStudy (anonymous):